# Group class operations and homological conditions

### v1

@@ 5. Groups satisfying Moore’s conjecture
Since $\mbox{Ext}^{n}_{kS^{\prime}}\!\left(M_{i},\mbox{res}_{S^{\prime}}^{S}N\right)\!=\mbox{Ext}^{n}_{kS_{i}}\!\left(\mbox{res}_{S_{i}}^{S^{\prime}}M,\mbox{res}_{S_{i}}^{S}N\right)=0$ for all $n>0$ and all $i$, the associated long exact sequence of Ext-groups that corresponds to the above short exact sequence of $kS^{\prime}$-modules implies that $\mbox{Ext}^{n}_{kS^{\prime}}\left(M,\mbox{res}_{S^{\prime}}^{S}N\right)=0$ for all $n>1$. This is the case for any $kS^{\prime}$-module $M$ and hence $\mbox{id}_{kS^{\prime}}\mbox{res}_{S^{\prime}}^{S}N\leq 1$. Then, Proposition 5.2(ii) implies that $\mbox{res}_{S^{\prime}}^{S}N\in{\tt Inj}(kS^{\prime})$.
We shall now prove that $N$ is an injective $kS$-module, using induction on the cardinality $\kappa$ of $S$. If $\kappa\leq\aleph_{0}$, then the result follows from the case considered above. If $\kappa$ is uncountable, then we may express $S$ as a continuous ascending union of subgroups $(S_{\lambda})_{\lambda<\kappa}$, each one having cardinality $<\kappa$. Our induction hypothesis implies that the $kS_{\lambda}$-module $\mbox{res}^{S}_{S_{\lambda}}N$ is injective for all $\lambda$. Let $M$ be a $kS$-module and define $M_{\lambda}=\mbox{ind}_{S_{\lambda}}^{S}\mbox{res}^{S}_{S_{\lambda}}M$ for all $\lambda$. We note that
@@
Here, $K$ is equal to the continuous ascending union of its submodules $(K_{\lambda})_{\lambda<\kappa}$. Since $K_{\lambda+1}/K_{\lambda}$ can be identified with the kernel of the (surjective) structure map $M_{\lambda}\longrightarrow M_{\lambda+1}$, we conclude that $\mbox{Ext}^{n}_{kS}(K_{\lambda+1}/K_{\lambda},N)=0$ for all $\lambda$ and $n>0$. Then, Eklof’s lemma [24, Theorem 7.3.4] implies that $\mbox{Ext}^{n}_{kS}(K,N)=0$ for all $n>0$; hence, the short exact sequence (3) above implies that $\mbox{Ext}^{n}_{kS}(M,N)=0$ for all $n>1$. Since this is the case for any $kS$-module $M$, we conclude that $\mbox{id}_{kS}N\leq 1$. Using once more Proposition 5.2(ii), we conclude that $N$ is injective. $\Box$
@@
As a consequence of Chouinard’s work [14], we know that finite groups satisfy Moore’s conjecture. In fact, all ${\scriptstyle{{\bf LH}}}\mathfrak{F}$-groups satisfy Moore’s conjecture; see [1] for the original (projective) version of the conjecture and [2] for the injective version of the conjecture (which implies the flat version, in view of [2, Proposition 2.6]). Let $\overline{\mathfrak{F}}$ be the smallest class of groups which contains the class $\mathfrak{F}$ of finite groups and is ${\scriptstyle{{\bf LH}}}$-closed, $\Phi$-closed and $\Phi_{flat}$-closed. We can describe $\overline{\mathfrak{F}}$-groups hierarchically, as explained in §$\S 1$.IV. The following Corollary is precisely Theorem B of the Introduction; it is itself an immediate consequence of Corollary 5.7(i), in view of the inclusion $\mathfrak{F}\subseteq\mathfrak{M}$.
@@
Acknowledgments. Wei Ren was supported by the National Natural Science Foundation of China (No. 11871125).

### v2

@@ 5. Groups satisfying Moore’s conjecture
Since $\mbox{Ext}^{n}_{kS^{\prime}}\!\left(M_{i},\mbox{res}_{S^{\prime}}^{S}N\right)\!=\mbox{Ext}^{n}_{kS_{i}}\!\left(\mbox{res}_{S_{i}}^{S^{\prime}}M,\mbox{res}_{S_{i}}^{S}N\right)=0$ for all $n>0$ and all $i$, the associated long exact sequence of Ext-groups that corresponds to the above short exact sequence of $kS^{\prime}$-modules implies that $\mbox{Ext}^{n}_{kS^{\prime}}\left(M,\mbox{res}_{S^{\prime}}^{S}N\right)=0$ for all $n>1$. This is the case for any $kS^{\prime}$-module $M$ and hence $\mbox{id}_{kS^{\prime}}\mbox{res}_{S^{\prime}}^{S}N\leq 1$. Since $H^{\prime}\cap S^{\prime}$ is a subgroup of finite index in $S^{\prime}$ and the restriction $\mbox{res}_{H^{\prime}\cap S^{\prime}}^{S^{\prime}}\mbox{res}_{S^{\prime}}^{S}N=\mbox{res}_{H^{\prime}\cap S^{\prime}}^{S}N=\mbox{res}_{H^{\prime}\cap S^{\prime}}^{H^{\prime}}\mbox{res}_{H^{\prime}}^{S}N$ is injective, as a restriction of the injective $kH^{\prime}$-module $\mbox{res}_{H^{\prime}}^{S}N$, Proposition 5.2(ii) implies that $\mbox{res}_{S^{\prime}}^{S}N\in{\tt Inj}(kS^{\prime})$.
We shall now prove that $N$ is an injective $kS$-module, using induction on the cardinality $\kappa$ of $S$. If $\kappa\leq\aleph_{0}$, then the result follows from the case considered above. If $\kappa$ is uncountable, then we may express $S$ as a continuous ascending union of subgroups $(S_{\lambda})_{\lambda<\kappa}$, each one having cardinality $<\kappa$. Our induction hypothesis implies that the $kS_{\lambda}$-module $\mbox{res}^{S}_{S_{\lambda}}N$ is injective for all $\lambda$. (Indeed, $\mbox{res}^{S}_{S_{\lambda}}N$ is injective when restricted to any finitely generated subgroup of $S_{\lambda}$, as well as to the subgroup $H^{\prime}\cap S_{\lambda}$, which has finite index in $S_{\lambda}$.) Let $M$ be a $kS$-module and define $M_{\lambda}=\mbox{ind}_{S_{\lambda}}^{S}\mbox{res}^{S}_{S_{\lambda}}M$ for all $\lambda$. We note that
@@
Here, $K$ is equal to the continuous ascending union of its submodules $(K_{\lambda})_{\lambda<\kappa}$. Since $K_{\lambda+1}/K_{\lambda}$ can be identified with the kernel of the (surjective) structure map $M_{\lambda}\longrightarrow M_{\lambda+1}$, we conclude that $\mbox{Ext}^{n}_{kS}(K_{\lambda+1}/K_{\lambda},N)=0$ for all $\lambda$ and $n>0$. Then, Eklof’s lemma [25, Theorem 7.3.4] implies that $\mbox{Ext}^{n}_{kS}(K,N)=0$ for all $n>0$; hence, the short exact sequence (4) above implies that $\mbox{Ext}^{n}_{kS}(M,N)=0$ for all $n>1$. Since this is the case for any $kS$-module $M$, we conclude that $\mbox{id}_{kS}N\leq 1$. Using once more Proposition 5.2(ii), we conclude that $N$ is injective. $\Box$
@@
As a consequence of Chouinard’s work [15], all finite groups satisfy Moore’s conjecture. This result was extended to the class of ${\scriptstyle{{\bf H}}}_{1}\mathfrak{F}$-groups in [1]. In fact, all ${\scriptstyle{{\bf LH}}}\mathfrak{F}$-groups satisfy Moore’s conjecture; see [2] for the original (projective) version of the conjecture and [3] for the injective version of the conjecture (which implies the flat version, in view of [3, Proposition 2.6]). Let $\overline{\mathfrak{F}}$ be the smallest class of groups which contains the class $\mathfrak{F}$ of finite groups and is ${\scriptstyle{{\bf LH}}}$-closed, $\Phi$-closed and $\Phi_{flat}$-closed; $\overline{\mathfrak{F}}$-groups can be described hierarchically, as we explained in §$\S 1$.IV. The following Corollary is precisely Theorem B in the Introduction. It is itself an immediate consequence of Corollary 5.7(i), in view of the inclusion $\mathfrak{F}\subseteq\mathfrak{M}$.
@@
Acknowledgments. It is a pleasure to thank Sergio Estrada, Gregory Kendall, Peter Kropholler and Olympia Talelli for useful comments and suggestions on an earlier draft of this paper. Wei Ren was supported by the National Natural Science Foundation of China (No. 11871125).